OpenStudy (anonymous):

By using the Ideal gas law variables and by using the variables m to represent mass and M to represent molar mass, derive an algebraic expression for this: the molar mass of a gas (without using "M" or "n" in your equation)

OpenStudy (aaronq):

PV=nRT n=m/M PV=mRT/M M=mRT/PV